OpenStudy (anonymous):

Steve drove at a constant rate to the beach for a vacation. In the equation below, t is the time in hours it took Steve to drive to the beach. 66t=330 What is the unit rate in the equation above? A. 5 miles per hour B. 264 miles per hour C. 66 miles per hour D. 330 miles per hour

OpenStudy (mathmale):

Josh: Please show how you arrive at your answers. In the long run you'd get better feedback by doing so.

OpenStudy (anonymous):

i only post the ones i have no clue on

OpenStudy (anonymous):

sorry i was on brb

OpenStudy (mathmale):

Are you sure that the original problem doesn't provide units of measurement with the "66" and "330?"

OpenStudy (anonymous):

no i'm saying i DON'T KNOW

OpenStudy (anonymous):

i'll check that out

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

not all of us are good at math xD

OpenStudy (mathmale):

One o the more common applications of algebra is the relationship Distance = Rate * Time. Here, "rate" refers to speed. d = r * t.

OpenStudy (mathmale):

Given d = r * t, solve for the rate, r.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

what is r 330?

OpenStudy (anonymous):

ok i see it's speed

OpenStudy (anonymous):

but i don't see where they give that is it the 330?

OpenStudy (mathmale):

Gave you an algebraic formula ( d = r*t ) and asked you to solve it algebraically for the rate (speed), r. Given d = rt, solve for r, please.

OpenStudy (mathmale):

This was your original post: Steve drove at a constant rate to the beach for a vacation. In the equation below, t is the time in hours it took Steve to drive to the beach. 66t=330

OpenStudy (mathmale):

Given d = r * t, solve for r.

OpenStudy (anonymous):

yes but if ir's r * t how do i find r?

OpenStudy (anonymous):

r anyway

OpenStudy (mathmale):

Solve for r algebraically. That means: Isolate r. If d = r * t, divide both sides of the equation by t. That will leave r alone on the right side.

OpenStudy (anonymous):

i still don't get it i'm right where i started

OpenStudy (mathmale):

d = r * t. distance equals rate times time. Solve for the rate, r, by dividing both sides of this equation by t: d r* t ---- = ----- t t Cancel out the "t" on the right side, please. Your result?

OpenStudy (faiqraees):

rate = t 330=66t since you have to make t the subject (t alone on lhs) divide 66 on both sides t = 330/66 t=.....?

OpenStudy (anonymous):

but if t is the distance then what's d?

OpenStudy (anonymous):

5

OpenStudy (mathmale):

Faiq: Joshua needs to solve for the RATE, not for the time.

OpenStudy (anonymous):

i have NO idea what you're talking about at all :l

OpenStudy (mathmale):

Have you not done any preparation? Read an introductory algebra book? It's expecting a lot, perhaps too much, if you come to OpenStudy to expect people to teach you concepts from the very beginning.

OpenStudy (anonymous):

forget it ima guess onto the next question for me

OpenStudy (faiqraees):

oh sorry t = speed correction